OpenStudy (yacoub1993):

Find the zeros of the polynomial function. f(x) = x3 + 5x2 - 4x - 20 A. x = -5, x = 4 B. x = -2, x = 2 C. x = -5, x = -2, x = 2 D. x = 5, x = -2, x = 2

hero (hero):

Hint: x^3 + 5x^2 = x^2(x + 5) -4x - 20 = -4(x + 5)

OpenStudy (yacoub1993):

B

hero (hero):

Well, that was definitely a guess.

OpenStudy (yacoub1993):

SORRY a

hero (hero):

Yet another guess. I suppose my hint was not helpful enough for you

hero (hero):

Trust me. Guessing is not the way.

OpenStudy (yacoub1993):

OK Explain Please

OpenStudy (yacoub1993):

ok

OpenStudy (yacoub1993):

wow it disappeared

hero (hero):

The best way to figure it out is to factor it. The best factoring method is "factor by grouping". With factor by grouping, you factor the first two terms x^2 + 5x^2, then then last two terms -4x - 20 When you factor each pair of terms you get: x^2(x + 5) and -4(x + 5) Notice that x + 5 is common to both factorizations. So then you factor out x + 5 to get (x + 5)(x^2 - 4) Now since x^2 - 4 is a difference of squares that factors to (x + 2)(x - 2) so the complete factorization of f(x) is f(x) = (x + 5)(x + 2)(x - 2) Then you set f(x) = 0 0 = (x + 5)(x + 2)(x - 2) And then set each factor equal to zero: 0 = x + 5 0 = x + 2 0 = x - 2 From there, you solve each for x.

OpenStudy (yacoub1993):

x=5 x=2 x=2

OpenStudy (yacoub1993):

x=-2 ast one

hero (hero):

You did not solve for x.

OpenStudy (yacoub1993):

i am confused

hero (hero):

I find it strange that I have to do every single step for you.

hero (hero):

0 = x + 5 Subtract 5 from both sides x = -5 0 = x + 2 Subtract 2 from both sides x = -2 0 = x - 2 Add 2 to both sides x = 2 So x = {-5, -2, 2}

OpenStudy (yacoub1993):

i did the same but the opposite way i subtracted 0 from both sides

OpenStudy (yacoub1993):

Thank you

hero (hero):

You have to ask yourself if it makes sense to subtract 0 from both sides. What will subtracting zero from both sides accomplish? I don't know if you are joking or not.